# Simultaneous Resonant and Broadband Detection of Ultralight Dark Matter and High-Frequency Gravitational Waves via Cavities and Circuits

### v2

@@ I Introduction
On the other hand, gravitational waves (GW) with frequencies higher than kHz can offer insights into early universe cosmology and particle physics beyond the standard model Aggarwal et al. (2021). Two noteworthy phenomena—the inverse Gertsenshtein effect Gertsenshtein (1962), akin to axion electrodynamics, and mechanical resonance Pegoraro et al. (1978a, b); Reece et al. (1984); Berlin et al. (2023)—have the capability to convert GW into photons in the presence of a background electromagnetic field. The largely unexplored parameter space of high-frequency GW (HFGW) has gained significant attention, leading to the implementation of axion haloscope experiments aimed at detecting its signatures Herman et al. (2021); Berlin et al. (2022a); Domcke et al. (2022); Sokolov (2022); Berlin et al. (2022b); Tobar et al. (2022); Berlin et al. (2023).
@@ IV.2 Physics Reach
Figure 2: Sensitivity reach of axion and dark photon DM, and HFGW, depicted with solid lines for single-mode and dashed lines for multi-mode detection limits. The sensitivity thresholds for single-mode are based on requiring Eqs. (44, 45, 46) to reach 11 11, while for multi-mode, the thresholds are derived from Eqs. (72, 73, 74, LABEL:eq:SRFEMA, LABEL:SNRMM). The integration time allocated for each $e$-fold of the target source frequency is $t_{e}=10^{7}$ s. Benchmark parameters for different experimental setups are discussed in Sec. IV.
@@ V Broadened Response Width in Multi-Mode Resonators
where the parameters $g$ and $G$ represent the couplings for beam-splitter-type and non-degenerate parametric interactions, respectively Russer and Russer (2011); Bergeal et al. (2010); Abdo et al. (2013); Chapman et al. (2016); Frattini et al. (2017). This system comprises $N+1$ modes denoted by $\hat{a}_{k}$, with each adjacent pair linked by the two types of interactions. The dynamics described by Eq. (49) can be interpreted as two copies of the Hatano-Nelson model, where two groups of quadratures are amplified in opposite directions McDonald and Clerk (2020). We designate $\hat{a}_{0}$ as the probing sensor, while the readout port is connected to the last mode $\hat{a}_{N}$, as illustrated in the top panel of Fig. 3. The application of the $N=1$ model of Eq. (49) for axion DM was previously discussed in Refs. Wurtz et al. (2021); Jiang et al. (2023).
@@ V.1 Scattering Matrices For Non-Hermitian Chain
Here, $\theta_{k}\equiv(\varphi^{G}_{k-1}-\varphi^{g}_{k-1}-\varphi^{G}_{k}-\varphi^{g}_{k})/2$. Consequently, the amplification process experiences a sequence of attenuation factors expressed as the product $\Pi_{k=1}^{N-1}\cos\theta_{k}$. It is important to observe that even with an initial calibration of all the relative phases, fluctuations in the phase of the pumping modes have the potential to contribute to $\varphi^{g/G}_{k}$ Wang et al. (2017).
@@ V.2 Scattering Matrices For Binary Tree
Figure 4: An illustration of a multi-mode resonator with a binary tree structure for $N=3$. Black lines indicate beam-splitter-like interactions, and blue dashed lines represent non-degenerate parametric interactions. At the lowest level, a network of $2^{N}$ sensors is positioned, facilitating the probing of $\Psi$. The remaining modes are auxiliary modes.
@@
In this setup, approximately $2^{N}$ sensors can be positioned at the bottom level to probe $\Psi$, each undergoing successive amplification as they propagate toward $\hat{b}$. The coherent nature of the source background can thus be leveraged to increase the sensitivity.
@@ V.3 Response Width in Multi-Mode Resonators
As discussed in Sec. IV, the response width governs the sensitivity reach of detectors. This width can be optimized numerically using Eq. (41), or approximated by the range in which intrinsic noise in $\hat{a}_{0}$ dominates over other noise contributions, as given by:
@@
To proceed, we first determine the range that satisfies a necessary condition from Eq. (61): the dominance of the left-hand side, representing intrinsic noise, over the readout noise on the right-hand side. This condition is characterized by a nearly constant PSD of approximately 11 11 for sufficiently small $\gamma$. By keeping $g$, $\gamma$, and $n_{\text{occ}}$ fixed, we then adjust $\gamma_{r}$ to match $\mathcal{J}$. According to Eq. (56), the condition $|S_{0r}|^{2}\gg 1$ holds within the region where $\Omega\ll\gamma_{r}$, but diminishes swiftly beyond this boundary. Leveraging this observation, we make the following substitutions: $\mathcal{J}$, $\gamma_{r}$, and $\Omega$ are substituted with $\Delta\omega_{r}$, leading to an optimized response width:
$\Delta\omega_{r}^{\rm opt}\simeq\Big{(}\gamma\,n_{\textrm{occ}}\,\mathcal{G}^{2N}\Big{)}^{1/(2N+1)},$ (62)
which converges to $2|g|$ for large $N$. Figure 5 depicts numerical instances of noise PSDs, with $\gamma_{r}=\Delta\omega_{r}^{\text{opt}}$ held constant and $\mathcal{J}$ varied. In these scenarios, the remaining condition from Eq. (61)—that intrinsic noise in $\hat{a}_{0}$ surpasses other intrinsic noise sources—is inherently fulfilled. The figure demonstrates that a smaller value of $\mathcal{J}$ mildly affects $\Delta\omega_{r}$, despite compressing the PSD within a narrower $\Omega$ range. By setting $\gamma_{r}=\mathcal{J}=\Delta\omega_{r}^{\text{opt}}$, we achieve a relatively flat PSD within $\Delta\omega_{r}$, rendering it robust against potential variations in $\mathcal{J}$, in light of the reasonable demands for dynamic range.
@@
Compared to single-mode resonators, the inclusion of both beam-splitter and non-degenerate parametric interactions with comparable interaction strengths $|g|\sim|G|$ effectively suppresses off-resonant phase shifts, as seen in the denominators of Eqs. (55, 60). This leads to a sequential broadening of the response width. Consequently, by extending the response width from an order of $n_{\text{occ}}\gamma$ to $2|g|$, multi-mode resonators significantly enhance the scan rate beyond that of single-mode resonators, as demonstrated in Fig. 5.
@@ VI Realization of Multi-Mode Resonators
One method to achieve the necessary couplings is through three-wave mixing, where one pumping mode’s frequency matches the sum or difference of the other two, thereby enabling non-degenerate parametric and beam-splitter-type interactions, respectively. Feasible implementations of three-wave mixing include the use of a Josephson Ring Modulator (JRM) Bergeal et al. (2010), Superconducting Nonlinear Asymmetric Inductive eLement (SNAIL) Frattini et al. (2017), and DC-driven Josephson junction effects Russer and Russer (2011). Notably, the JRM structure has recently been utilized to demonstrate response width broadening in a $N=1$ resonator composed of two circuit modes Jiang et al. (2023). Thus, we use JRMs as the benchmark for constructing multi-mode resonators, as illustrated in the bottom panel of Fig. 3.
@@ VI.1 Multi-Mode Resonators Connected by Josephson Ring Modulators
The use of a JRM to connect two modes with a third pumping mode is illustrated in the bottom panel of Fig. 3. Each JRM consists of four Josephson junctions, each with the same critical current $I_{0}$, labeled by $i=1,2,3,4$. This results in a total Hamiltonian given by Bergeal et al. (2010):
@@ VI.2 Potential Challenges
The implementation of three-wave mixing, as discussed above, has been previously demonstrated in Ref. Bergeal et al. (2010). Additionally, a prototype that realizes both $g$ and $G$ types of couplings simultaneously using a Josephson Ring Modulator (JRM) has shown the potential to amplify a mimicking axion signal, achieving a 5.6-fold enhancement in SNR2 with $N=1$ Jiang et al. (2023). However, constructing a multi-mode resonator capable of significantly broadening the response bandwidth presents several challenges, which we will briefly address along with potential solutions below.
@@
  • Another potential challenge is the phase fluctuation of the pump mode $\Phi_{p}$, which deviates from the frequencies at $\omega_{\rm rf}^{k-1}\pm\omega_{\rm rf}^{k}$. In the non-Hermitian chain model discussed in Sec. V.1, resultant phase fluctuations $\varphi^{g/G}_{k}$ can quench the amplification Wang et al. (2017). Additionally, misalignment of these phases can suppress the amplification of quadratures as described in Eq. (58). However, the binary tree model in Sec. V.2 completely circumvents this issue, as both quadratures are equally amplified.
  • A large noise occupation number $n_{\rm occ}$ can challenge the assumption $\Phi_{k}\ll\varphi_{0}$, as $\Phi_{k}\sim\sqrt{n_{\rm occ}}\kappa_{k}$. According to Eqs. (12) and (13), zero-point uncertainties $\kappa_{k}$ are adjustable by modifying the structure of the readout antenna in a cavity or manipulating the inductance and capacitance of an LC circuit. However, reducing $\kappa_{k}$ excessively will diminish the magnitudes of $|g|$ and $|G|$ as defined in Eq. (69), necessitating a higher Josephson energy $E_{J}$ to compensate.

### v3

@@ I Introduction
On the other hand, gravitational waves (GW) with frequencies higher than kHz can offer insights into early universe cosmology and particle physics beyond the standard model Aggarwal et al. (2021, 2025). Two noteworthy phenomena—the inverse Gertsenshtein effect Gertsenshtein (1962), akin to axion electrodynamics, and mechanical resonance Pegoraro et al. (1978a, b); Reece et al. (1984); Berlin et al. (2023)—have the capability to convert GW into photons in the presence of a background electromagnetic field. The largely unexplored parameter space of high-frequency GW (HFGW) has gained significant attention, leading to the implementation of axion haloscope experiments aimed at detecting its signatures Herman et al. (2021); Berlin et al. (2022a); Domcke et al. (2022); Sokolov (2022); Berlin et al. (2022b); Tobar et al. (2022); Berlin et al. (2023); Navarro et al. (2024).
@@ IV.2 Physics Reach
Figure 2: Sensitivity reach of axion and dark photon DM, and HFGW, depicted with solid lines for single-mode and dashed lines for multi-mode detection limits. The sensitivity thresholds for single-mode are based on requiring Eqs. (44, 45, 46) to reach 11 11, while for multi-mode, the thresholds are derived from Eqs. (80, 81, 82, LABEL:eq:SRFEMA, LABEL:SNRMM). The integration time allocated for each $e$-fold of the target source frequency is $t_{e}=10^{7}$ s. Benchmark parameters for different experimental setups are discussed in Sec. IV.
@@ V Broadened Response Width in Multi-Mode Resonators
where the parameters $g$ and $G$ represent the couplings for beam-splitter-type and non-degenerate parametric interactions, respectively Russer and Russer (2011); Bergeal et al. (2010a); Abdo et al. (2013); Chapman et al. (2016); Frattini et al. (2017). This system comprises $N+1$ modes denoted by $\hat{a}_{k}$, with each adjacent pair linked by the two types of interactions. The dynamics described by Eq. (49) can be interpreted as two copies of the Hatano-Nelson model, where two groups of quadratures are amplified in opposite directions McDonald and Clerk (2020). We designate $\hat{a}_{0}$ as the probing sensor, while the readout port is connected to the last mode $\hat{a}_{N}$, as illustrated in the top panel of Fig. 3. The application of the $N=1$ model of Eq. (49) for axion DM was previously discussed in Refs. Wurtz et al. (2021); Jiang et al. (2023).
@@ V.1 Scattering Matrices For Non-Hermitian Chain
Here, $\theta_{k}\equiv(\varphi^{G}_{k-1}-\varphi^{g}_{k-1}-\varphi^{G}_{k}-\varphi^{g}_{k})/2$. Consequently, the amplification process experiences a sequence of attenuation factors expressed as the product $\Pi_{k=1}^{N-1}\cos\theta_{k}$. It is important to note that even with an initial calibration of all relative phases, fluctuations in the phases of the pumping modes can still contribute to $\varphi^{g/G}_{k}$ Wang et al. (2017). Therefore, it is necessary to recalibrate the phase before each data acquisition cycle, as we discuss later.
@@ V.2 Scattering Matrices For Binary Tree
Figure 4: An illustration of a multi-mode resonator with a binary tree structure for $N=3$. Black lines indicate beam-splitter-like interactions, and blue dashed lines represent non-degenerate parametric interactions. At the lowest level, a network of $2^{N}$ sensors can be positioned, facilitating the probing of $\Psi$. The remaining modes are auxiliary modes.
@@
In this setup, approximately $2^{N}$ sensors can be positioned at the bottom level to probe $\Psi$, each undergoing successive amplification as they propagate toward $\hat{b}$. The coherent nature of the background source can be exploited to enhance sensitivity: signals from different probing sensors can be coherently combined at the readout port, boosting the signal PSD by a factor of $2^{2N}$. Although intrinsic thermal noise from each probing sensor also contributes, these fluctuations are uncorrelated across modes and thus grow only as $2^{N}$. This results in a net enhancement of the SNR2 by a factor of $2^{N}$.
@@ V.3 Response Width in Multi-Mode Resonators
As discussed in Sec. IV, the response width determines the sensitivity reach of detectors. This width can be optimized numerically using Eq (41) or approximated by identifying the range where intrinsic noise from $\hat{a}_{0}$ dominates over other noise contributions, as expressed in:
@@
Here, the left-hand side represents intrinsic noise from $\hat{a}_{0}$, while the right-hand side accounts for readout noise and intrinsic noise from higher resonators.
The goal of optimization is to identify the best set of free parameters that maximizes the response width, as specified by Eq. (61), for a fixed number of chains, $N$. In principle, this condition depends on five parameters: $|g|$, $\gamma$, $n_{\textrm{occ}}$, $\gamma_{r}$, and $\mathcal{J}$. We assume that $\gamma$ and $n_{\textrm{occ}}$ are optimized experimentally to be as small as possible, and that $|g|$ is maximized to enhance $|S_{0r}|$. The remaining free parameters, $\gamma_{r}$ and $\mathcal{J}$, are tuned for optimization.
To begin, we examine the properties on the left-hand side of Eq. (61), particularly the scattering matrix $|S_{0r}|^{2}$. From Eqs. (56) and (57), we obtain
$|S_{0r}(\Omega=\Delta\omega_{r})|^{2}$ (62)
$=$ $\,\frac{4\gamma\gamma_{r}\mathcal{G}^{2N}|f_{0}|^{2}}{|\gamma_{r}f_{N}+f_{N+1}|^{2}}$
$\approx$  $\,4\gamma\gamma_{r}\mathcal{G}^{2N}/\left|\mathcal{J}^{N}\gamma_{r}\sum_{j=0}^{\lfloor N/2\rfloor}C_{N-j}^{j}\left(-\mathrm{i}\frac{\Delta\omega_{r}}{\mathcal{J}}\right)^{N-2j}\right.$
$\left.+\,\mathcal{J}^{N+1}\sum_{j=0}^{\lfloor(N+1)/2\rfloor}C_{N+1-j}^{j}\left(-\mathrm{i}\frac{\Delta\omega_{r}}{\mathcal{J}}\right)^{N+1-2j}\right|^{2}$
$\approx$ $\,\frac{4\gamma\gamma_{r}\mathcal{G}^{2N}}{\Delta\omega_{r}^{2N}(\gamma_{r}+\Delta\omega_{r})^{2}}.$
In the third line, we neglect $\gamma$ in the expression for $f_{N}=\Sigma_{j=0}^{\lfloor N/2\rfloor}C_{N-j}^{j}(\gamma-\mathrm{i}\Omega)^{N-2j}\mathcal{J}^{2j}$, taking $f_{0}=1$, and evaluate the bandwidth at the boundary of the response width, $\Omega=\Delta\omega_{r}$. The last line is derived from the following three regions:
  • When $\mathcal{J}<\Delta\omega_{r}$, the highest-order polynomial term with $j=0$ dominates in each sum, leading directly to the final expression.
  • When $\mathcal{J}>\Delta\omega_{r}$, the condition is inconsistent. In this region, $|S_{0r}|^{2}$ remains nearly constant for $\Omega<\mathcal{J}$, dominated by the 00 0-th order term in one of the polynomial series with either $j=\lfloor N/2\rfloor$ or $\lfloor(N+1)/2\rfloor$. $|S_{0r}|^{2}$ only begins to decrease significantly for $\Omega>\mathcal{J}$, ensuring that the bandwidth satisfies $\Delta\omega_{r}>\mathcal{J}$.
  • When $\mathcal{J}\approx\Delta\omega_{r}$, both sums become approximately 1 based on numerical evaluation. Consequently, the term $|\cdots|^{2}$ can be approximated as $|\mathcal{J}^{N}\gamma_{r}+\mathcal{J}^{N+1}|^{2}$, which leads to the final expression by setting $\mathcal{J}\approx\Delta\omega_{r}$.
@@
To proceed, we first determine the range of parameters that satisfies the necessary condition from Eq. (61), where the dominance of the left-hand side over the readout noise—characterized by a nearly constant PSD of approximately 11 11 for sufficiently small $\gamma$—is required. This condition is expressed as
$|S_{0r}(\Omega=\Delta\omega_{r})|^{2}n_{\rm occ}\approx\frac{4n_{\rm occ}\gamma\gamma_{r}\mathcal{G}^{2N}}{\Delta\omega_{r}^{2N}(\gamma_{r}+\Delta\omega_{r})^{2}}\geq 1.$ (63)
which must be satisfied to achieve an optimal response width. Maximizing $\Delta\omega_{r}$ for this relation is achieved by tuning $\gamma_{r}$, which leads to the simple conclusion that $\gamma_{r}\approx\Delta\omega_{r}$. It is important to note that this relation depends only marginally on the value of $\mathcal{J}$. As we will discuss later, variations in $\mathcal{J}$ affect the flatness of the PSD within $\Delta\omega_{r}$, since in the region where $\Omega\gg\mathcal{J}$, the PSD decreases rapidly.
Given this, we take $\mathcal{J}=\gamma_{r}=\Delta\omega_{r}$ as our benchmark parameters. This yields the expression $|S_{0r}|^{2}\sim\gamma\mathcal{G}^{2N}/\Delta\omega_{r}^{2N+1}$, and the optimized response width, derived from Eq. (63), becomes
$\Delta\omega_{r}^{\rm opt}\simeq\Big{(}\gamma\,n_{\textrm{occ}}\,\mathcal{G}^{2N}\Big{)}^{1/(2N+1)}.$ (64)
In the limit of large $N$, this width converges to $\mathcal{G}$, provided that $\gamma n_{\rm occ}$ remains finite. Moreover, we can use $\mathcal{J}=\Delta\omega_{r}^{\rm opt}$ to solve for the exact values of $\mathcal{J}$ and $\mathcal{G}$, which satisfy
$\mathcal{J}^{\rm opt}=\left(\gamma n_{\rm occ}(\sqrt{|g|^{2}-\left(\mathcal{J}^{\rm opt}\right)^{2}}+|g|)^{2N}\right)^{1/(2N+1)}.$ (65)
For small $N$, the ratio $\mathcal{G}/|g|$ is close to 22 22, but it approaches 11 11 as $N$ increases. Meanwhile, $\Delta\omega_{r}=\mathcal{J}$ approaches $|g|$ as well.
@@
Figure 5 shows numerical examples of noise PSDs, with $\gamma_{r}=\Delta\omega_{r}^{\text{opt}}$ held constant and $\mathcal{J}$ varied. In these cases, the condition from Eq. (61)—that the intrinsic noise in $\hat{a}_{0}$ (solid lines) surpasses other intrinsic noise sources (dashed lines)—is inherently satisfied. The figure illustrates that a smaller value of $\mathcal{J}$ has a minimal effect on $\Delta\omega_{r}$, even though it compresses the PSD within a narrower $\Omega$ range, thus validating our earlier argument. By setting $\gamma_{r}=\mathcal{J}=\Delta\omega_{r}^{\text{opt}}$, we achieve a relatively flat PSD within $\Delta\omega_{r}$, making it robust against potential variations in $\mathcal{J}$, especially considering the reasonable demands for the experimental dynamic range.
For varying $\mathcal{J}$, the PSD amplitude in the flat region can be derived as:
$|S_{0r}(\Omega=0)|^{2}n_{\rm occ}\approx 4n_{\textrm{occ}}\gamma\gamma_{r}\mathcal{G}^{2N}/(\gamma_{r}\mathcal{J}^{N}+\mathcal{J}^{N+1})^{2}.$ (66)
which scales as $\propto\mathcal{J}^{-2N}$ when $\mathcal{J}\ll\gamma_{r}$. The PSD remains flat within $\Omega\ll\mathcal{J}$ and rapidly decreases outside of this range, as shown by the green line in Fig. 5.
Figure 6: Contributions to the noise PSDs from different mode levels for the $N=5$ case shown in Fig. 5. The solid lines represent the even-level contributions, which exhibit a nearly flat spectrum within the bandwidth, while the dashed lines correspond to the odd-level contributions, which scale as $\propto\Omega^{2}/\gamma_{r}^{2}$. The amplitudes of each class follow a geometric sequence with a ratio of $(\gamma_{r}/\mathcal{G})^{4}$, with the dominant contribution coming from the $N=0$ mode.
In Fig. 6, we show the contribution to PSDs from different levels for the $N=5$ case presented in Fig. 5. Here, the relatively flat PSDs come from even-level modes, while the power-law dependent PSDs, scaling as $\propto\Omega^{2}/\gamma_{r}^{2}$, arise from odd-level modes. The amplitudes of each class follow a geometric sequence with a ratio of $(\gamma_{r}/\mathcal{G})^{4}$, with the dominant contribution from the $N=0$ mode. Within the response width, their PSDs can be approximated as:
$|S_{kr}|^{2}n_{\textrm{occ}}=\frac{\gamma_{r}^{2k}}{\mathcal{G}^{2k}}\times\left\{\begin{aligned} 4,\qquad\quad&\text{for }N\in\text{ even}\\ \,(k+1)^{2}\frac{\Omega^{2}}{\gamma_{r}^{2}}.\quad&\text{for }N\in\text{ odd}\end{aligned}\right.$ (67)
The total contribution from odd-level modes is negligible compared to the even ones.
Higher-order contributions to noise can be accounted for by introducing an effective thermal occupation number, $n_{\rm occ}^{\rm eff}$, where the ratio $n_{\rm occ}^{\rm eff}/n_{\rm occ}$ represents the total noise PSD amplitude relative to the one from the 00 0-th mode only. This is given by:
$|S_{0r}|^{2}n_{\rm occ}^{\rm eff}=\sum_{i=0}^{N}|S_{ir}|^{2}n_{\textrm{occ}}.$ (68)
Using the geometric series with a ratio of $(\gamma_{r}/\mathcal{G})^{4}$ from the even-level modes, we obtain:
$n_{\rm occ}^{\rm eff}=n_{\textrm{occ}}\frac{\mathcal{G}^{4}-(\gamma_{r}/\mathcal{G})^{4[N/2]}\gamma_{r}^{4}}{\mathcal{G}^{4}-\gamma_{r}^{4}}.$ (69)
These higher-order contributions can limit the optimal number of levels. According to Eq. (43), the sensitivity now becomes $\text{SNR}^{2}(\overline{\omega}_{\Psi})\propto\Delta\omega_{r}^{\text{opt}}/(n_{\rm occ}^{\textrm{eff}})^{2}$. Since both $\Delta\omega_{r}^{\rm opt}$ and $n_{\rm occ}^{\rm eff}$ increase with $N$, one can numerically solve for the optimal $N$ under a given $|g|/(\gamma n_{\rm occ})$. The optimal value of $N$ is always an odd number, which provides a larger response width $\Delta\omega_{r}^{\text{opt}}$ compared to the next smaller even number, with $n_{\rm occ}^{\rm eff}$ remaining nearly unchanged. A higher value of $g/(\gamma n_{\textrm{occ}})$ typically leads to a higher optimal number of levels, as the ratio of the geometric series becomes suppressed, i.e., $(\gamma_{r}/\mathcal{G})\propto(\gamma n_{\rm occ}/g)^{4}$.
For the examples shown in Figs. 5 and 6 with $g/\gamma=10^{6}$ and $n_{\rm occ}=10$, the optimal level is $N=9$, while for SRF detection with $g/\gamma=10^{12}$ and $n_{\rm occ}=100$, the optimal number is $N=17$. At these optimal levels, the response width for both cases is greater than $80\%$ of the width when $N\rightarrow\infty$.
@@
Historically, the goal of achieving a broad bandwidth with a large response originated in the context of gravitational wave detection, particularly in the concept of the white-light cavity Wicht et al. (1997), where an anomalous medium with negative dispersion was proposed to counteract phase shifts in the off-resonant region. In our case, the combined use of beam-splitter and non-degenerate parametric couplings functions analogously: the response is first significantly enhanced by the non-degenerate parametric coupling $|G|$, while the addition of beam-splitter interaction with comparable strength $|g|\sim|G|$ flattens the response by reducing the peak, as seen in the denominators of Eqs. (55, 60). This mechanism leads to a sequential broadening of the response width as the signal propagates. Consequently, by extending the response width from an order of $n_{\text{occ}}\gamma$ to $2|g|$, multi-mode resonators dramatically enhance the scan rate beyond that of single-mode resonators, as illustrated in Fig. 5.
@@ VI Realization of Multi-Mode Resonators
One method to achieve the necessary couplings is through three-wave mixing, where one pumping mode’s frequency matches the sum or difference of the other two, thereby enabling non-degenerate parametric and beam-splitter-type interactions, respectively. Feasible implementations of three-wave mixing include the use of a Josephson Ring Modulator (JRM) Bergeal et al. (2010a), Superconducting Nonlinear Asymmetric Inductive eLement (SNAIL) Frattini et al. (2017), and DC-driven Josephson junction effects Russer and Russer (2011). Notably, the JRM structure has recently been utilized to demonstrate response width broadening in a $N=1$ resonator composed of two circuit modes Jiang et al. (2023). Thus, we use JRMs as the benchmark for constructing multi-mode resonators, as illustrated in the bottom panel of Fig. 3.
@@ VI.1 Multi-Mode Resonators Connected by Josephson Ring Modulators
The use of a JRM to connect two modes with a third pumping mode is illustrated in the bottom panel of Fig. 3. Each JRM consists of four Josephson junctions, each with the same critical current $I_{0}$, labeled by $i=1,2,3,4$. This results in a total Hamiltonian given by Bergeal et al. (2010a):
@@ VI.2 Potential Challenges
The implementation of three-wave mixing, as discussed above, has been previously demonstrated in Ref. Bergeal et al. (2010a). Additionally, a prototype that realizes both $g$ and $G$ types of couplings simultaneously using a Josephson Ring Modulator (JRM) has shown the potential to amplify a mimicking axion signal, achieving a 5.6-fold enhancement in SNR2 with $N=1$ Jiang et al. (2023). However, constructing a multi-mode resonator capable of significantly broadening the response bandwidth presents several challenges, which we will briefly address along with potential solutions below.
@@
  • Similarly, the operating temperature of the JRM must be kept below a certain threshold, typically ensuring that the thermal occupation number is dominated by quantum fluctuations. In the benchmark examples shown in Figs. 5 and 6, we assume $n_{\rm occ}=10$ for illustration. However, the environmental temperature surrounding the JRMs may differ from that inside the resonators. For instance, SRF cavities are typically cooled to around $1.8\,\mathrm{K}$ using liquid helium immersion Tang et al. (2024). Given that the setup is considerably larger than the cavity itself, there is enough space to accommodate an additional small dilution refrigerator, which can provide the mK-level temperatures required for JRM operation. Alternatively, the SRF cavity could be placed in the K-level stage of a sufficiently large dilution refrigerator to provide adequate cooling power, while the JRM is positioned in the mK stage of the same refrigerator to ensure proper operation.
  • Another potential challenge is the phase fluctuation of the pump mode $\Phi_{p}$, which can deviate from the frequencies at $\omega_{\rm rf}^{k-1}\pm\omega_{\rm rf}^{k}$. In the non-Hermitian chain model discussed in Sec. V.1, the resulting phase fluctuations, denoted $\varphi^{g/G}_{k}$, can quench the amplification Wang et al. (2017). Misalignment of these phases can suppress the amplification of the quadratures, as described in Eq. (58).
    For the non-Hermitian chain structure, phase misalignment can be recalibrated Jiang et al. (2023). The phase of the JRM can be adjusted by manipulating the phases of the drive currents that generate the beam-splitter-type and non-degenerate parametric interactions (i.e., the pump current $I_{g/G}$). Since phase alignment maximizes the output power, each JRM’s phase can be tuned during calibration to optimize this. This power maximization criterion can make the calibration process automatic.
    Typically, the JRM phase remains stable for much longer than the duration of the typical data acquisition cycle. The total integration time can be divided into several cycles, with calibration of all relevant parameters performed before data collection in each cycle.
    On the other hand, the binary tree model in Sec. V.2 completely circumvents this issue, as both quadratures are equally amplified.
@@
  • Another source of nonlinearities arises from the expansion in Eq. (72), which requires that $\Phi_{k}\ll\varphi_{0}$ to prevent nonlinear effects from JRMs from invalidating our derivation. As shown in Eq. (11), $\Phi_{k}$ can be approximated by $\Phi_{k}\sim\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle^{1/2}\kappa_{k}$. The steady-state energy in the resonators, from $\hat{a}_{0}$ to $\hat{a}_{N}$, grows geometrically with the ratio $(\mathcal{G}/\gamma_{r})^{2}>1$. Therefore, we only need to ensure that the energy in the last mode, $\hat{a}_{N}$, satisfies $\Phi_{N}\ll\varphi_{0}$.
    The steady-state energy of $\hat{a}_{N}$ can be estimated as:
    $\langle\hat{a}_{N}^{\dagger}\hat{a}_{N}\rangle\approx\frac{1}{\gamma_{r}}\Delta\omega_{r}\,|S_{0r}(\Omega=0)|^{2}n_{\rm occ}\approx 4,$ (78)
    where we assume that $|S_{0r}|^{2}$ is roughly flat within the response bandwidth to simplify the expression, and we use the benchmark response width from Eq. (64) in the final estimate.
    Meanwhile, from Eqs. (12) and (13), we see that zero-point fluctuations $\kappa_{k}$ can be adjusted by modifying the readout antenna structure in a cavity or by manipulating the inductance and capacitance of an LC circuit. This allows $\kappa_{k}$ to be reduced, ensuring that $\Phi_{k}\ll\varphi_{0}\approx 1.65$ in natural units. However, excessively reducing $\kappa_{k}$ would decrease the magnitudes of $|g|$ and $|G|$ (as defined in Eq. (76)), necessitating a higher Josephson energy $E_{J}$ to compensate.
  • JRMs may experience power saturation when the injected power is too large. Since the power injected from the resonator $\hat{a}_{k}$ into the JRM is approximately given by $g\,\omega_{\rm rf}^{k}\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle$, and the coupling strength $g$ between each resonator remains nearly constant, it suffices to ensure that the JRM connected to $\hat{a}_{N}$ does not reach saturation.
    In the benchmark case, with $\langle\hat{a}_{N}^{\dagger}\hat{a}_{N}\rangle\approx 4$ (from Eq. (78)), and considering $\omega_{\rm rf}^{N}\approx 2\pi\,\mathrm{GHz}$ and $g\approx\omega_{\rm rf}^{N}$, the maximum injected power is approximately $g\,\omega_{\rm rf}^{N}\langle\hat{a}_{N}^{\dagger}\hat{a}_{N}\rangle\approx-107.8\,\mathrm{dBm}$. According to Ref. Bergeal et al. (2010b), the saturation power of the JRM can reach around $-105\,\mathrm{dBm}$. Therefore, in the benchmark case considered in this study, the JRMs will not experience power saturation.
@@
  • The readout system is connected only to the final, highest-level mode in the network. This connection—typically implemented via an antenna or transmission line—is essentially identical to the standard readout approach used in single-mode resonators. However, a key distinction in the multi-mode setup is the significantly broadened response width, which requires the amplification chain to perform reliably across a wide frequency range. This imposes practical constraints on the choice of amplifiers. In particular, traveling-wave parametric amplifiers are well-suited for this application, as they offer broadband operation while maintaining quantum-limited noise performance.
  • In the binary tree configuration, achieving the full coherent enhancement of the signal—potentially improving the SNR2 by a factor of $2^{N}$—requires phase matching across all probing sensors at the lowest level. Several factors can affect the relative phases of signals arriving at the readout port. First, the intrinsic phase coherence of the background source plays a key role. For ultralight DM, the coherence length is typically $\sim 10^{3}$ times longer than the Compton wavelength, which is itself comparable to or larger than the detector scale. This ensures phase coherence across the sensor array. A similar argument applies to HFGW detection using SRF cavities or LC circuits, where the GW wavelength is much longer than the detector dimensions. In the case of heterodyne upconversion detection, the background electromagnetic field used for mixing introduces an additional phase contribution. Furthermore, spatial separations between sensors can introduce relative phase shifts.
    A viable calibration strategy is to inject a known coherent signal into all probing sensors simultaneously and adjust the relative phases by tuning the lengths of superconducting delay lines, thereby maximizing the resulting signal at the readout port.